• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 00:14:16 on July24,2005

    Entry number 149247

    keyword=Swing Shift Summary

    Swing Shift Summary
    -------------------------
    Continue x,y scan in counting mode to study rates vs position.  We are using 
    0.4 uA of beam and a 3x3 raster on LHe for these tests.  Each run consists of 
    1M events.
    
    Run      BPMA         BPMB
    2237   (1.6,0.8)   (-1.4,-1.1)
    2238   (1.55,0.8)  (-1.4,+1.0)
    2239   (1.57,0.07) (-1.4,+1.9)
    2240   (1.5,0.05)  (-1.4,2.8) *only 300k events
    2241   (1.6,0.2)   (-1.4,2.8)
    
    Now, start X scans:
    
    2242   (1.4,0.2)   (-3.1,0.05)
    2243   (1.7,-0.06) (3.1,-0.2)
    2244   (1.5,-0.1)  (1.0,-0.1)
    2245   (1.5,-0.1)  (-1.1,-0.04)
    
    18:35  - go back to nominal beam positions
             BPMA (1.6,0); BPMB (-1.4,0) +/-0.2
    
    18:36 - change Septum current to +10% 
            SepR = 275 A (these are readback current, not setpoints)
    	SepL = 280 A
    
    18:40 - Asked for no beam so that Kent could test dithering control 
    signals/software.
    
    18:47 - Kent done.  Ask for beam back.
    
    Had to wait a few minutes to get bpm positions right.
    
    18:55 - Run 2246 started.  Septum at +10%
    
    19:02 - change septum current to +20%
            SepR = 299 A
    	SepL = 304 A
    
    19:07 - Run 2247 started.  Septum at +20%.
    
    19:20 - Switch over to HAPPEX DAQ from Spectrometer DAQ following checklist.
            Also waiting for septum to come back to nominal values.
    
    19:32 -  Ask for 7x7 (MCC = 14x14) raster.
             Verified with spot++.
    
    19:47 - Start HAPPEX run 3787 and set Detector and Blumi HV for 4 uA running.
    
    Begin Current/fan speed scans:
    
    19:59 - Run 3788: fan speed = 60 Hz, 4 uA, end of run has no beam
    
    20:07 - Run 3789: fan speed = 40.5 Hz, 4 uA, beginning of run has no beam.
    
    Lowered HV in preparation for higher beam current.
    
    20:16 - Ask MCC for 10 uA.
    
    20:18 - Set Det and blumi voltages (Run 3790)
    
    20:24 - Run 3791: fan speed = 40.5 Hz, 10 uA, end of run has no beam
    
    20:30 - Change fan spped to 60 Hz.
    
    20:31 - Run 3792: fan speed = 60 Hz, 10 uA, beginning of run has no beam
    
    20:39 - Lower HV on dets and blumis to productions values.
    
    20:41 - Ask MCC for 30 uA
    
    20:43 - Run 3793 - set HV for 30 uA
    
    20:51 - Run 3794: fan speed = 60 Hz, 30 uA, end of run has no beam
    
    20:57 - Change fan speed to 40 Hz.
    
    20:58 - Run 3795: fan speed = 40.5 Hz, 30 uA, beginning of run has no beam.
    
    21:08 - Set fan speed back to 60 Hz, set all HV back to nominal, and ask for 35 
    uA beam current to begin production-style running.
    
    21:16 - Run 3796: production run, 35 uA, 7x7 raster, 60 Hz fan speed.
    
    22:12 - Run 3797: production run.
    
    22:45 - Right Q2 quenched.  See halog:14924
    5
           
    22:53 - Called MCC for controlled access to go down and reset Q2.
    
    23:10 - Bryan and Lisa went down to the hall to try to reset Q2.
    
            Reset of magnet didn't work - tried several times.  Called RC, he said 
    it was ok to call tech-on-call.
    
    00:10 - Todd is on his way in.